# Wannier states and spin supersolid physics in the triangular antiferromagnet K2Co(SeO3)2

### v1

@@ III Model and theory
Since the Hilbert space of the effective low-energy model is the Wannier subspace ${\cal W}$, it is first necessary to introduce a suitable algebra of operators that act on this highly constrained space. This is achieved by introducing dimer coverings on the dual honeycomb lattice. As shown in Figure 7(a) and (b), a dimer is placed on the dual link perpendicular to each frustrated bond (a bond connecting two parallel spins) of a given spin configuration. We denote a specific dimer covering by ${\cal C}$. Because we are focusing on only one of the two states related by time-reversal symmetry, the proposed duality transformation results in a 2:1 mapping.
@@
Figure 7: a) Triangular lattice (dashed line) and the associated honeycomb dual lattice (bold line). The green bonds denote the “special” ones used in the unitary transformation. b) A resonant hexagon, the red boxes represent the bonds covered by a dimer. c) Effect of the Hamiltonian over a resonant hexagon, denoted by the blue box, and the number of special bonds covered by dimers $N_{s}$. d) Representation of the calculation of $\langle S^{z}_{i}S^{z}_{j}\rangle$, for $ij$ connected by the blue dashed line. Since the line crosses for an odd number (1) of non-covered bonds in the dual lattice the correlator takes a negative value, as is expected by inspection.
@@
In this language, ${\cal H}_{\rm eff}$ is simply a quantum dimer model with a ring-exchange term that acts on each pair of adjacent hexagons on the dual honeycomb lattice (corresponding to each bond $\langle i,j\rangle$ of the triangular lattice).
@@
In other words, ${\cal H}_{\rm eff}$ connects basis states that present a double-hexagon resonance as sketched in Figure 7(c). The ground state is then obtained by optimizing the energy gain produced by this dimer resonance.
@@
with $\innerproduct{{\cal C^{\prime}}}{{\cal C}}=\delta_{\cal C^{\prime}C}$ and $\sum_{\cal C}|\phi({\cal C})|^{2}=1$. If a unitary transformation $\hat{U}$ exists that changes the sign of ${\cal H}_{\rm eff}$, the low-energy model can be mapped to the one that is for negative $J_{xy}$, i.e., non-frustrated XY interaction. Consider the lattice depicted in Figure 7(a), where specific bonds—designated as “special”—are marked with green lines. For any given dimer covering $\cal C$, we define $N_{s}({\cal C})$ as the number of special bonds occupied by dimers. As illustrated in Figure 7(c), a resonance within any resonant hexagon changes $N_{s}({\cal C})$ by ±$\pm 2$. To capture this behavior, we introduce the following unitary transformation on the dimer basis [50]:
$|\mathcal{C}^{\prime}\rangle:=\hat{U}|\mathcal{C}\rangle=i^{N_{s}(\mathcal{C})}|\mathcal{C}\rangle$ (7)
@@
where $i$ is the imaginary unit. Consequently, basis elements connected by ${\cal H}_{\rm eff}$ differ in their $N_{s}$ values by integer multiples of ±$\pm 2$. Defining $|\Psi_{\alpha}^{\prime}\rangle=\hat{U}|\Psi_{\alpha}\rangle$, we obtain
$\matrixelement{\Psi_{\gamma}^{\prime}}{{\cal H}_{\rm eff}}{\Psi_{\alpha}^{\prime}}$ =$=\sum_{\{\cal C\}}~{}\!\!\!\!-i^{N_{s}({\cal C^{\prime}})}i^{N_{s}({\cal C})}\phi_{\gamma}^{*}({\cal C^{\prime}})\phi_{\alpha}({\cal C})\matrixelement{\cal C^{\prime}}{{\cal H}_{\rm eff}}{\cal C}$
=$=\sum_{\{\cal C\}}~{}i^{\pm 2}\phi_{\gamma}^{*}({\cal C^{\prime}})\phi_{\alpha}({\cal C})\matrixelement{\cal C^{\prime}}{{\cal H}_{\rm eff}}{\cal C}$
=$=\sum_{\{\cal C\}}~{}-\phi_{\gamma}^{*}({\cal C^{\prime}})\phi_{\alpha}({\cal C})\matrixelement{\cal C^{\prime}}{{\cal H}_{\rm eff}}{\cal C}$
$=\matrixelement{\Psi_{\gamma}}{-{\cal H}_{\rm eff}}{\Psi_{\alpha}},$ (8)
implying that $\hat{U}$ maps ${\cal H}_{\rm eff}$ into $-{\cal H}_{\rm eff}$ without affecting the structure of the wave function. Since the unitary transformation preserves the energy spectrum, thermodynamic properties, which depend solely on the energy eigenvalues, remain unchanged.
Consider the spin-spin correlation function $\langle S^{z}_{i}S^{z}_{j}\rangle$. As illustrated in Figure 7(d), a line connecting sites $i$ and $j$ intersects a set of honeycomb lattice edges $\langle ab\rangle$. The value of $S^{z}_{i}S^{z}_{j}$ depends on the parity of unoccupied edges along this line: it equals $+S^{2}$ when the number of intersected edges without dimers is even, and $-S^{2}$ when this number is odd. Now, define the dimer number operator $\hat{n}_{ab}:=(1+4S^{z}_{i}S^{z}_{j})/2$ where $ab$ refers to the honeycomb lattice edge dual to the triangular lattice edge $ij$. Using this operator, the correlation function becomes a non-local string operator, expressed as:
@@
$\langle S^{z}_{i}S^{z}_{j}\rangle=\frac{1}{4}\Big{\langle}\!\!\!\!\!\prod_{\begin{subarray}{c}\langle ab\rangle\\ ~{}\text{cut by}~{}ij\end{subarray}}\!\!\!\!\!\left(2\hat{n}_{ab}-1\right)\Big{\rangle}.$ (9)
The product has $2^{|j-i|}$ terms (with $|j-i|$ being the distance between $j$ and $i$ measured in units of the triangular lattice constant). These terms can be evaluated using the Grassmannian integral method. Because these operators are diagonal in the dimer basis, $\matrixelement{\Psi^{\prime}}{S^{z}_{i}(t)S^{z}_{j}(0)}{\Psi^{\prime}}=\matrixelement{\Psi}{S^{z}_{i}(t)S^{z}_{j}(0)}{\Psi}$. This, however, does not hold for $\langle S^{x}_{i}(t)S^{x}_{j}(0)\rangle$ or $\langle S^{y}_{i}(t)S^{y}_{j}(0)\rangle$ correlation functions, which take a more complex form after the transformation.
@@
Our approach to solving the Hamiltonian ${\cal H}$ is illustrated in Figure 8. To address the sign problem, we project the Hamiltonian onto the Wannier space and employ the transformation introduced by Wang. This transformation effectively flips the sign of the $xy$-term in the Hamiltonian without altering the thermodynamic properties or the $\langle S_{i}^{z}(t)S_{j}^{z}(0)\rangle$ correlation functions. While implementing an efficient QMC algorithm under the constraints of ${\cal W}$ implied a significant computational challenge, we bypass this difficulty by solving the non-frustrated XXZ Hamiltonian and neglecting second-order corrections in $J_{xy}$. The validity of this approximation is demonstrated by the results discussed in the following sections.
IV Model validation
As discussed in the previous section, the current model predicts three BKT transitions at zero field. Each BKT transition is expected to produce a broad peak in the specific heat, consistent with the sequence of three peaks observed in our heat capacity data shown in Figure 4. The two upper BKT transition temperatures, $T_{{\rm BKT},1}$ and $T_{{\rm BKT},2}$, are on the order of 10 K and 1 K, respectively, with the lowest transition at $T_{{\rm BKT},3}\approx 0.35~{}\textrm{K}$. Although the system is technically three-dimensional, and the upper two transitions are replaced by a single phase transition at $T_{c1}\simeq T_{{\rm BKT},1}$, the corresponding anomaly in the specific heat is too weak to be detectable. As a result, the specific heat curve is well approximated by the 2D model. These results are also in good agreement with the finite-temperature Lanczos results presented in [51] for the same model. The weaker onset of XY ordering at $T_{c2}\simeq T_{{\rm BKT},3}\approx 0.35~{}\textrm{K}$ is more clearly observed in Figure 9, where a slight elbow marks the transition.
To further validate the proposed spin Hamiltonian and the low-energy model derived in the previous section, we compute the specific heat curve $C_{v}(T,H)$ in the low-temperature regime for the different values of the magnetic field $\mu_{0}H$, as well as the uniform magnetization curve $M(H)$, up to the values corresponding to the onset of the $1/3$ magnetization plateau.
As shown in Figure 3, the calculated magnetization curve is in good agreement with the experimental data, predicting a critical field $\mu_{0}H_{c}\approx 0.85~{}\textrm{T}$, which matches well with the observed value. Although the theoretical calculation shows a slightly non-linear increase of $M$ with $H$, which is characteristic of 2D systems but seems to be absent in the measured curve. We note that LSWT theory predicts a slightly lower critical field $\mu_{0}H_{c}=0.71$ T.
@@ IV Model validation
The specific heat $C_{v}=R~{}\partial_{T}\langle\mathcal{H}\rangle$ is computed for $\mu_{0}H=0,0.25$ and $0.5$ T and temperatures in the interval $0.16~{}\text{K}<T<3.27~{}\text{K}$. The lowest temperature reached by QMC calculations is slightly higher than the experimental value ($T=0.12$ K). This limitation, combined with finite-size effects, likely explains the minor discrepancies observed in comparisons with experimental data.
The comparison with the experiment is presented in Figure 9. The phonon and Schottky contributions were subtracted from the measured data as in Figure 4(b). The above-mentioned low-temperature broad peaks are successfully reproduced by the QMC simulations of the low-energy effective model confirming the validity of the Hamiltonian parameters of K2Co(SeO3)2. Moreover, as shown in Figure 10, the low-energy model also reproduces the temperature dependence of the magnetic entropy as it approaches the Wannier value.
@@
Figure 9: Magnetic specific heat of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ for different fields $\mu_{0}H$ = 0, 0.25 and 0.5 T applied along the $c$ axis. Circles: experimental data from Ref. [28] with phonon and nuclear contribution subtracted as described in the text. The line is a guide for the eye. Squares: QMC calculations (this work). The black arrows indicate the BEC transition.
@@
V Quantum Spin Dynamics and Inelastic Neutron Scattering
In this section, we present and analyze high-resolution, low-energy inelastic neutron scattering data collected in the spin supersolid phase under an applied magnetic field. We first examine the experimental neutron scattering results, then compare these findings with the QMC simulations, and conclude by presenting and discussing the high-energy excitations at zero magnetic fields.
In Figure 5 we show the low-energy magnetic excitation spectrum of K2Co(SeO3)2 measured by high-resolution inelastic neutron scattering at AMATERAS as a function of energy and high-symmetry directions in the reciprocal space at $T=70~{}\textrm{mK}$ and zero magnetic field. The incident neutron energy is $E_{i}=1.69$ meV giving a resolution of $23\mu$eV at elastic scattering (to be compared to the resolution $\Delta E\approx$ 36 $\mu$eV for the data reported in [28]). To enhance the counting statistics of the neutron intensity, the data has been averaged over equivalent paths in the reciprocal space accessible in the experiment, as shown in Figure 20. The spectrum is dominated by the broad, gapless continuum of excitations, deviating drastically from the predictions of the linear spin wave theory (magnon dispersion in solid lines), and no sharp, coherent magnon modes can be resolved even with the improved energy resolution.
This is more obvious in the energy-dependent cuts plotted in Figure 11 at several representative wave vectors: M point [Figure 11(b)], half way between the K and M points [Figure 11(c)], and half way between the M and $\Gamma$ points [Figure 11(e)]. A sharp excitation feature may be present at the lower boundary of the continuum for these wave vectors, whereas it can not be resolved at the wave vector halfway between the $\Gamma$ and K point [Figure 11(d)]. Note that the intensity of the scattered neutrons is plotted in the logarithmic scale, and the corresponding plots with linear scale are displayed in the insets. The observation of a sharp feature at the lower boundary of a continuum of excitations in low-dimensional quantum magnets is often interpreted as a bound state of fractionalized spin excitations [52]. In the present study, we cannot resolve any gap between the sharp excitation and the rest of the continuum within our experimental resolution, which seems to support such a scenario. Nevertheless, one can never exclude the possibility that there exist closely spaced coherent magnon modes where the difference in energy is much smaller than our experimental resolution. Even then, our high-resolution INS data imposed strong constraints on the upper limit of the energy spacing between potential discrete magnon modes. Further, our data allow us to make an accurate estimate on the spectral weight contributions of the sharp magnon modes, if existing, at the lower boundary of the continuum. At the wave vectors investigated [Figure 11(b),(c),(e)], the weight of the sharp excitation (represented by a shaded Gaussian in which FWHM corresponds to the energy resolution) is only about 40%, implying that at least 60% of the total spectral weight is contributed by the continuum.
At the K point, in addition to the gapless Goldstone excitations as expected from the spontaneously broken U(1) symmetry in the supersolid phase, a gapped excitation is observed at approximately $E\approx 0.06$ meV [Figure 11(a)]. This excitation corresponds to the previously described pseudo-Goldstone mode and is prominently visible in the energy-momentum plot of the magnetic excitation spectrum near the K point, as illustrated in the inset of Figure 5.
Figure 11: Energy cuts through the INS data shown in Figure 5 at representative wave vectors as indicated. Both longitudinal and transverse spin fluctuations contribute to scattering, the intensity being proportional to $S^{zz}(\mathbf{q},\omega)+S^{\bot}(\mathbf{q},\omega)$. (a) K point, (b) M point, (c) half way between K and M, (d) half way between $\Gamma$ and K, and (e) half way between M and $\Gamma$. The solid lines represent QMC calculations for $S^{zz}(\mathbf{q},\omega)$ only. An arbitrary global scale factor is applied to match the intensity of the M point. The measured spectra have been averaged over equivalent wave vectors accessible in the experiment. Note the y-axis is in logarithmic scale. Insets show the same plot with y-axis in linear scale. Shaded Gaussians represent the energy resolution. The dashed lines are fits to a constant background and the elastic incoherent scattering peak.
To better understand the nature of the continuum of excitations in the supersolid phase of K2Co(SeO3)2, we have investigated their evolution in a magnetic field applied along the magnetic easy axis (i.e. $c$ axis) at LET. In Figure 12 (a)-(d), we show the false-color plots of the energy-momentum cuts of the low-energy magnetic excitation spectra measured at $T$ = 50 mK and in a field of $\mu_{0}H$ = 0, 0.25, 0.5, and 0.75 T, respectively. The spectra were plotted after subtracting the background scattering from the sample environment measured at $\mu_{0}H$ = 7 T (note there is no magnetic excitation below 1 meV at 7 T). As before, to enhance the counting statistics, the spectra presented have been averaged over equivalent paths in the reciprocal space as illustrated in Figure 21. Typical averaged energy-dependent cuts of the spectra at the K and M points are shown in Figure 13. The energy resolution of this dataset ($\Delta E=40$ $\mu$eV) is lower than that obtained at AMATERAS [Figure 5 and Figure 11]. While the low-energy spectrum at zero magnetic field [Figure 12(a) and Figure 13(a)] appears qualitatively similar in both measurements, the application of a magnetic field drastically reconstructs the excitation spectrum. To emphasize the distinctions from LSWT, Figure 18 illustrates the dynamical spin structure factor (DSSF) calculated using this approach.
@@ V Quantum Spin Dynamics and Inelastic Neutron Scattering
The roton-like dispersion minimum of the zero-field gapless branch [Figure 12(a)] is progressively suppressed with increasing magnetic field [Figure 12(b)], consistent with the system becoming more semi-classical as it approaches the 1/3 magnetization plateau. In contrast, the dip at the M point persists for the gapped branch [Figure 12(b)]. Another indication of the system’s evolution towards a semi-classical regime is the field-induced suppression of the zero-field dip in the lower boundary of the continuum, located halfway between the $\Gamma$ and K points [see Figure 12(a) and Figure 5]. This dip, absent in the spin-wave theory predictions shown by solid lines, is completely suppressed at $\mu_{0}H=0.25~{}\textrm{T}$ [Figure 12(b)]. Consistent with this observation, Figure 12(b) also demonstrates that the excitations become sharper and more intense as they approach the $\Gamma$ point. While LSWT accurately reproduces the slope of the excitations in the $E\rightarrow 0$ limit at both the $\Gamma$ and K points, notable discrepancies arise between the measured and predicted excitations at higher energies.
At $\mu_{0}H=0.5$ T, the magnetic excitation spectrum undergoes further reconstruction, as shown in Figure 12(c). The gapped excitation at the K point shifts to a higher energy of 0.26 meV [see also Figure 13(c)]. Most remarkably, the continuum of excitations moves towards higher energy with increasing magnetic field, and appears to be associated with the gapped excitation branch. Moreover, the degeneracy at the M point is completely lifted, with the energy of the excitation in the gapped branch increasing to 0.35 meV [Figure 13(g)]. The roton-like dip in the gapped branch becomes hardly visible [Figure 12(c)]. Overall, the excitation intensity becomes much weaker at this field (note the $\times$ 2 factor in color scale).
Finally, the continuum of excitations disappears at $\mu_{0}H=0.75$ T, just below the quantum phase transition to the uud plateau phase at $\mu_{0}H_{c}\approx 0.8$ T, as shown in Figure 12(d). The excitations evolve into sharp magnons, resembling those observed in the plateau phase at $\mu_{0}H_{c}=1.5$ T, as previously reported [28]. It is worth noting that, as discussed in the previous section, LSWT predicts a lower critical field implying that the calculated magnon bands in Figure12(d) are gapped. This reaffirms that the continuum of excitations in the supersolid state of K2Co(SeO3)2 is linked to the “superfluidity” of the transverse $xy$-moment in the Y-structure.
While the gapless branch is well described by LSWT, the gapped branch exhibits minor deviations. These discrepancies arise because LSWT is not fully accurate near the quantum critical point. Although the spin structure closely resembles the collinear uud configuration, quantum fluctuations significantly reduce the ordered moment from its saturated value [28]. Furthermore, the scattering appears diffusive: the energy width of the excitations slightly exceeds the resolution [Figure 13(h)], and weak scattering is observed at the K point near 0.5 meV, just below the gapped excitation branch [see Figure 12(d) and Figure 13(d)].
Figure 12: False-color plots of neutron scattering intensity measured on K2Co(SeO3)2 on the LET spectrometer at $T$ = 50 mK as a function of energy and wave vector along high-symmetry directions of the Brillouin zone in a magnetic field applied along the $c$ axis: (a) $\mu_{0}H=0$, (b) 0.25 T, (c) 0.5 T, and (d) 0.75 T. The incident neutron energy is $E_{i}$ = 2.15 meV, the resolution is 0.04 meV at the elastic line. The data were integrated over ±$\pm 0.75$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Fig. 21. The spectra measured at 7 T have been subtracted to remove the background from the sample environment. Note that in (d) the residual scattering at the K point near 0.4 meV is due to spurious scattering from the cryomagnet and imperfect background subtraction. Solid lines are the dispersion curves calculated by the linear spin wave theory. The white hexagon represents the Brillouin zone boundary. $\Gamma$, $K$ and $M$ label high symmetry points in the reciprocal space.
Figure 13: Magnetic field evolution of the neutron scattering intensity in K2Co(SeO3)2, as measured at the LET spectrometer with $E_{i}=2.15$ meV, as a function of energy transfer at the (a-d) K point and (e-h) M point. A magnetic field of $\mu_{0}H$ = 0, 0.25, 0.5, and 0.75 T, respectively, is applied along the crystallographic $c$ axis. The spectra have been averaged over equivalent wave vectors accessible in the experiment. The spectra measured at 7 T have been subtracted to remove the background from the sample environment. Note in (d) the scattering near 0.4 meV is spurious and due to imperfect backbround subtraction. The scale on the y-axis is linear. The shaded Gaussians represent the calculated energy resolution. The solid lines represent QMC calculations for $S^{zz}(\mathbf{q},\omega)$ only. An arbitrary global scale factor is applied to match the intensity of the M point at $H$ = 0. The dashed lines are fits to the elastic peak.
In the previous section, we validated our numerical approach by calculating the heat capacity and total magnetization and comparing the results with experimental data. We demonstrated that this methodology performs exceptionally well in accurately describing these quantities. The low-energy excited states, with energies on the order of $E\sim J_{xy}$, reside within the Wannier space, and it is anticipated that our approach can effectively capture these excitations as well. However, a significant limitation of our methodology is our inability to compute the ${\cal S}^{xx}$ and ${\cal S}^{yy}$ correlation functions. Consequently, we focus on the ${\cal S}^{zz}$ component of the DSSF, defined as
@@
with $S^{z}_{{\bf q}}=\frac{1}{{\cal N}}\sum_{j}e^{-{\bf q}\cdot{\bf r}_{j}}S^{z}_{j}$ with $j$ running over all lattice sites. Although a direct comparison with experimental data is not possible —since the data includes all three components of the dynamical spin structure factor— several features observed experimentally are clearly reflected in our simulations, underscoring the strong quantum origin of the magnetic excitations.
[…]
@@
Figure 14: ${\cal S}^{zz}$ component of the dynamical spin structure factor computed with QMC and analytic continuation for (a) $\mu_{0}H=0$ T, (b) $\mu_{0}H=0.25$ T and (c) $\mu_{0}H=0.5$ T. The white dots are the first poles of ${\cal S}^{zz}$ extracted with the method presented in Appendix  B.
@@
The primary features of the experimental spectrum are well reproduced at zero field [Figure 14(a)]. Our simulations successfully capture the roton-like modes, and the continuum along the $\Gamma$K path shows strong agreement with the experimental data. In Figure 11, we present cuts at different momenta for the zero-field calculations, emphasizing the accurate prediction of both the potential quasiparticle peak and the broad continuum. To ensure consistency, the intensities are rescaled to match the M point. However, a significant discrepancy with the experimental data is the absence of an extended continuum at the M point in the simulations. We conjecture that this continuum primarily originates from the ${\cal S}^{xx}({\bf q},\omega)$ and ${\cal S}^{yy}({\bf q},\omega)$ components of the dynamical spin structure factor, a hypothesis supported by the Lanczos results reported in [53].

### v2

@@ II.5 Low energy theory
Since the Hilbert space of the effective low-energy model is the Wannier subspace ${\cal W}$, it is first necessary to introduce a suitable algebra of operators that act on this highly constrained space. This is achieved by introducing dimer coverings on the dual honeycomb lattice. As shown in Figure 14(a), a dimer is placed on the dual link perpendicular to each frustrated bond (a bond connecting two parallel spins) of a given spin configuration. We denote a specific dimer covering by ${\cal C}$. Because we are focusing on only one of the two states related by time-reversal symmetry, the proposed duality transformation results in a 2:1 mapping.
@@
Figure 14: Dual lattice and dimer model. a) Triangular lattice (grey line) and the associated honeycomb dual lattice (bold line). The green bonds denote the “special” ones used in the unitary transformation. The blue dashed lines show a resonant hexagon and the red boxes represent the dimer covering. b) Effect of the effective Hamiltonian over a resonant hexagon, the same that in Figure (a), the number of special bonds covered by dimers is represented by $N_{s}$.
@@
In this language, ${\cal H}_{\rm eff}$ is simply a quantum dimer model with a ring-exchange term that acts on each pair of adjacent hexagons on the dual honeycomb lattice (corresponding to each bond $\langle i,j\rangle$ of the triangular one).
@@
In other words, ${\cal H}_{\rm eff}$ connects basis states that present a double-hexagon resonance as sketched in Figure 14(b). The ground state is then obtained by optimizing the energy gain produced by this dimer resonance.
@@
with $\innerproduct{{\cal C}_{\nu}}{{\cal C}_{\mu}}=\delta_{\nu\mu}$ and $\sum_{\mu}|\psi_{\mu}|^{2}=1$. If a unitary transformation $\hat{U}$ exists that reverses the sign of ${\cal H}_{\rm eff}$, the low-energy model can be mapped onto one with a negative $J_{xy}$, corresponding to a non-frustrated XY interaction. Consider the lattice depicted in Figure 14(a), where specific bonds—designated as “special”—are marked with green lines. For any given dimer covering $\cal C$, we define $N_{s}({\cal C})$ as the number of special bonds occupied by dimers. As illustrated in Figure 14(b), a resonance within any resonant hexagon changes $N_{s}({\cal C})$ by ±$\pm 2$. To capture this behavior, we introduce the following unitary transformation on the dimer basis [29]:
|$\ket{\mathcal{C}^{\prime}}:=\hat{U}\ket{\mathcal{C}}=e^{i\frac{\pi}{2}{N_{s}(\mathcal{C})}}~{}\ket{\mathcal{C}}$ (7)
@@
where $i$ is the imaginary unit. Consequently, basis elements connected by ${\cal H}_{\rm eff}$ differ in their $N_{s}$ values by integer multiples of ±$\pm 2$. Defining |$\ket{\Psi^{\prime}}=\hat{U}\ket{\Psi}$ and |$\ket{\Phi^{\prime}}=\hat{U}\ket{\Phi}$, we obtain
@@
$\begin{split}\langle{\Phi}^{\prime}|\mathcal{H}_{\mathrm{eff}}|{\Psi}^{\prime}\rangle&=\sum\limits_{\mu\nu}e^{i\frac{\pi}{2}\left[N_{s}(\mathcal{C}_{\mu})-N_{s}(\mathcal{C}_{\nu})\right]}\,\phi_{\nu}^{*}\psi_{\mu}\langle\mathcal{C}_{\nu}|\mathcal{H}_{\mathrm{eff}}|\mathcal{C}_{\mu}\rangle\\[6.0pt] &=\sum\limits_{\mu\nu}e^{\pm i\pi}\,\phi_{\nu}^{*}\psi_{\mu}\langle\mathcal{C}_{\nu}|\mathcal{H}_{\mathrm{eff}}|\mathcal{C}_{\mu}\rangle\\[6.0pt] &=\sum\limits_{\mu\nu}-\phi_{\nu}^{*}\psi_{\mu}\langle\mathcal{C}_{\nu}|\mathcal{H}_{\mathrm{eff}}|\mathcal{C}_{\mu}\rangle\\[6.0pt] &=\langle\Phi|-\mathcal{H}_{\mathrm{eff}}|\Psi\rangle\end{split}$ (8)
implying that $\hat{U}$ maps ${\cal H}_{\rm eff}$ into $-{\cal H}_{\rm eff}$. Since the unitary transformation preserves the energy spectrum, thermodynamic properties, which depend solely on the energy eigenvalues, remain unchanged. The same holds for observables that are diagonal in the dimer basis, implying that $\langle\Psi^{\prime}|S^{z}_{i}(t)S^{z}_{j}(0)|\Psi^{\prime}\rangle=\langle\Psi|S^{z}_{i}(t)S^{z}_{j}(0)|\Psi\rangle$. However, this does not apply to correlators like $\langle S^{x}_{i}(t)S^{x}_{j}(0)\rangle$ or $\langle S^{y}_{i}(t)S^{y}_{j}(0)\rangle$, which take a more complex form after the transformation.
@@
We summarize our strategy for solving the Hamiltonian ${\cal H}$ in Figure 15. To address the sign problem, the original XXZ Hamiltonian is projected onto the Wannier space where the transformation introduced by Wang permits reverse the sign of the $xy$-term without altering the thermodynamic properties or the $\langle S_{i}^{z}(t)S_{j}^{z}(0)\rangle$ correlation functions. While implementing an efficient QMC algorithm under the constraints of ${\cal W}$ implied a significant computational challenge, we bypass this difficulty by solving the non-frustrated XXZ Hamiltonian and neglecting second-order corrections in $J_{xy}$. The validity of this approximation is demonstrated by the results discussed in the following sections.
II.6 QMC simulations
To further validate the proposed spin Hamiltonian and the low-energy model derived in the previous section, we computed the uniform magnetization curve $M(\mu_{0}H)$, up to the values corresponding to the onset of the $1/3$ magnetization plateau. As shown in Figure 12, the calculated magnetization curve is in good agreement with the experimental data, predicting a critical field $\mu_{0}H_{c}\approx 0.85~{}\textrm{T}$, which matches well with the observed value. The theoretical calculation shows a slightly non-linear increase of $M$ with $H$, which is characteristic of 2D systems but seems to be absent in the measured curve. Small discrepancies with the experimental data are likely due to the fact that the current mapping is exact only in the Ising limit and at zero field. As a result, small higher-order corrections in $J_{xy}$ are expected to introduce discrepancies with the experimental results. While we do not expect a perfect match, we anticipate a good agreement, given the small ratio $\alpha$ and the low field, which makes it plausible that the main contributions come from the Wannier states.
@@ II.6 QMC simulations
The specific heat $C_{v}=R~{}\partial_{T}\langle\mathcal{H}\rangle$ is computed for $\mu_{0}H=0,0.25$ and $0.5$ T and the low temperature interval $0.16~{}\text{K}<T<3.27~{}\text{K}$. The comparison with the experiment is presented in Figure 16. The phonon and Schottky contributions were subtracted from the measured data as in Figure 3(b). The mentioned low-temperature broad peaks are successfully reproduced by the QMC simulations of the low-energy effective model, confirming the validity of the Hamiltonian parameters of K2Co(SeO3)2. The lowest temperature reached by QMC calculations is slightly higher than the experimental value ($T=0.12$ K). This limitation, combined with finite-size effects, likely explains the minor discrepancies observed in comparisons with experimental data.
Moreover, as shown in Figure 17, the low-energy model also reproduces the temperature dependence of the magnetic entropy as it approaches the Wannier value. These results are also in good correspondence with previous Lanczos ones  [53].
@@
Figure 16: Experimental and calculated magnetic specific heat of K2Co(SeO3)2. Magnetic specific heat of $\text{K}_{2}\text{Co}(\text{Se}\text{O}_{3})_{2}$ for different fields $\mu_{0}H$ = 0, 0.25 and 0.5 T applied along the $c$ axis. Circles: experimental data from Ref. [33] with phonon and nuclear contribution subtracted as described in the text. The line is a guide for the eye. Squares: QMC calculations (this work). The black arrows indicate the BEC transition.
@@
The low-energy excited states, with energies on the order of $E\sim J_{xy}$, reside within the Wannier space, and it is anticipated that our approach can effectively capture these excitations as well. However, a significant limitation of our methodology is our inability to compute the ${\cal S}^{xx}$ and ${\cal S}^{yy}$ correlation functions. Consequently, we focus on the ${\cal S}^{zz}$ component of the DSSF, defined as
@@
where $S^{z}_{{\bf q}}=\frac{1}{\sqrt{{\cal N}}}\sum_{j}e^{-{\bf q}\cdot{\bf r}_{j}}S^{z}_{j}$, ${\cal N}$ is the number of sites and $j$ runs over all lattice sites.
Although a direct comparison with experimental data is not possible—since it includes all three components of the dynamical spin structure factor—the strong anisotropy of the $g$-tensor, discussed in Section II.3, indicated that the calculation should account for most of the measured spectral weight. Several experimentally observed features are clearly reflected in our simulations, highlighting the strong quantum nature of the magnetic excitations.
@@
Figure 18: Evolution of the ${\cal S}^{zz}$ component of the DSSF in the supersolid phase in a magnetic field, obtained by QMC calculations. ${\cal S}^{zz}$ component of the dynamical spin structure factor computed with QMC and analytic continuation for (a) $\mu_{0}H=0$ T, (b) $\mu_{0}H=0.25$ T and (c) $\mu_{0}H=0.5$ T. The white dots are the first poles of ${\cal S}^{zz}$ extracted with the method presented in Supplementary Note 5.
@@
The key features of the experimental spectrum are well reproduced at zero field [Figure 18(a)]. Our simulations accurately capture the roton-like minima, and the continuum along the $\Gamma$–K path shows strong agreement with the experimental data. In Figure 6, we plot the calculated momentum-resolved cuts (red solid lines), highlighting the precise prediction of both the potential quasiparticle peak and the broad continuum. To ensure consistency, intensities are rescaled to match the M point. Additionally, the bottom edge of the continuum is well described by the first pole of the QMC data [Figure 5, 10(a)], with a perfect match in the excitation velocity near K. A significant discrepancy with the experimental data is the absence of an extended continuum at the M point in the simulations. We conjecture that this continuum primarily originates from the ${\cal S}^{xx}({\bf q},\omega)$ and ${\cal S}^{yy}({\bf q},\omega)$ components of the dynamical spin structure factor, a hypothesis supported by the Lanczos results reported in [54].